# Single-minus gluon tree amplitudes are nonzero

### v1

@@ Abstract
In general, n2n-2 is actually not the maximally allowed number of plus gluons. In this paper, we show that n1n-1 plus (or “single-minus”) amplitudes are in fact allowed [^2] Witten [5] notes that single-minus amplitudes are supported at a point in twistor space; see also [6]. with restricted “half-collinear” kinematics. [^3] The half-collinear condition can be viewed as restricting the ingoing and outgoing momenta to a one-dimensional null circle on the celestial torus at the boundary of Klein space. The amplitude is divided into chambers whose walls are regions where sums of various subsets of the half-collinear momenta are orthogonal as described below. The (stripped) amplitudes are piecewise-constant integers in each chamber. The assignment of its value to each chamber is determined from the perturbative Berends–Giele recursion [15], which is equivalent to Feynman diagrams.
@@
where $z_{ij}\equiv z_{i}-z_{j}$ and $\tilde{z}_{ij}\equiv\tilde{z}_{i}-\tilde{z}_{j}$. We take our polarization vectors to be
@@ I.2 The recursion relation
$V_{\tilde{\lambda}_{1}\cdots\tilde{\lambda}_{n}}$ $=\prod_{k=1}^{n-1}\operatorname{sg}_{k,k+1}\Theta\mathopen{}\mathclose{{\left(-\frac{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{1\cdots k}\tilde{\lambda}_{k+1\cdots m}}}\right]}{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{k}\tilde{\lambda}_{k+1}}}\right]}}}\right),$ (20)
@@
where the ordered partition now has $A\geq 1$ parts, while
@@ II.3.3 Evaluating $\bar{V}_{\tilde{\lambda}_{2}\cdots\tilde{\lambda}_{n}}$
$\bar{V}_{\tilde{\lambda}_{2}\cdots\tilde{\lambda}_{n}}=\prod_{n=2}^{n}\operatorname{sg}_{m,m+1}\Theta\mathopen{}\mathclose{{\left(\frac{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{2\cdots m}\tilde{\lambda}_{m+1\cdots n}}}\right]}{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{m}\tilde{\lambda}_{m+1}}}\right]}}}\right).$ (52)

### v2

@@ Abstract
In general, n2n-2 is actually not the maximally allowed number of plus gluons. In this paper, we show that n1n-1 plus (or “single-minus”) amplitudes are in fact allowed even at tree level [^2] Witten [5] notes that single-minus tree amplitudes are supported at a point in twistor space; see also [6]. All-plus and single-minus amplitudes are generically corrected at loop level [15, 16]. with restricted “half-collinear” kinematics. [^3] The half-collinear condition can be viewed as restricting the ingoing and outgoing momenta to a one-dimensional null circle on the celestial torus at the boundary of Klein space. The amplitude is divided into chambers whose walls are regions where sums of various subsets of the half-collinear momenta are orthogonal as described below. The (stripped) amplitudes are piecewise-constant integers in each chamber. These values are determined from the perturbative Berends–Giele recursion [17], which is equivalent to Feynman diagrams.
@@
where $z_{ij}:=z_{i}-z_{j}$ and $\tilde{z}_{ij}:=\tilde{z}_{i}-\tilde{z}_{j}$. We take our polarization vectors to be
@@ I.2 The recursion relation
$V_{\tilde{\lambda}_{1}\cdots\tilde{\lambda}_{n}}$ $=\prod_{k=1}^{n-1}\operatorname{sg}_{k,k+1}\Theta\mathopen{}\mathclose{{\left(-\frac{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{1\cdots k}\tilde{\lambda}_{k+1\cdots n}}}\right]}{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{k}\tilde{\lambda}_{k+1}}}\right]}}}\right),$ (20)
@@
where the ordered partition of $(2\ldots n)$ now has $A\geq 1$ parts, while
@@ II.3.3 Evaluating $\bar{V}_{\tilde{\lambda}_{2}\cdots\tilde{\lambda}_{n}}$
$\bar{V}_{\tilde{\lambda}_{2}\cdots\tilde{\lambda}_{n}}=\prod_{m=2}^{n-1}\operatorname{sg}_{m,m+1}\Theta\mathopen{}\mathclose{{\left(\frac{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{2\cdots m}\tilde{\lambda}_{m+1\cdots n}}}\right]}{\mathopen{}\mathclose{{\left[\tilde{\lambda}_{m}\tilde{\lambda}_{m+1}}}\right]}}}\right).$ (52)